solve by the elimination method
5x+5y=-7
7x-3y=17
The solution is?

Answer by njusa(12)

5x+5y=-7... a
7x-3y=17... b
15x+15y=-21 ... 3*a
35x-15y=85 .... 5*b
50x = 64 .... a+b
x = 32/25
substitute this back into a and you will get the y. check by subtituting both x and y into b.